### v3

# VersusDebias: Universal Zero-Shot Debiasing for Text-to-Image Models via SLM-Based Prompt Engineering and Generative Adversary

Abstract
With the rapid development of Text-to-Image (T2I) models, biases in human image generation against demographic social groups become a significant concern, impacting fairness and ethical standards in AI. Some researchers propose their methods to tackle with the issue. However, existing methods are designed for specific models with fixed prompts, limiting their adaptability to the fast-evolving models and diverse practical scenarios. Moreover, they neglect the impact of hallucinations, leading to discrepancies between expected and actual results. To address these issues, we introduce VersusDebias, a novel and universal debiasing framework for biases in arbitrary T2I models, consisting of an array generation (AG) module and an image generation (IG) module. The self-adaptive AG module generates specialized attribute arrays to post-process hallucinations and debias multiple attributes simultaneously. The IG module employs a small language model to modify prompts according to the arrays and drives the T2I model to generate debiased images, enabling zero-shot debiasing. Extensive experiments demonstrate VersusDebias’s capability to debias any models across gender, race, and age simultaneously. In both zero-shot and few-shot scenarios, VersusDebias outperforms existing methods, showcasing its exceptional utility. Our work is accessible at https://github.com/VersusDebias/VersusDebias to ensure reproducibility and facilitate further research.
1 Introduction
Text-to-Image (T2I) models, as a crucial part of Artificial Intelligence Generated Content (AIGC) technology, are evolving at an extraordinary speed driven by open-source initiatives by companies such as Stability AI (Rombach et al. 2022; Luo et al. 2023; Betker et al. 2023), enabling even non-experts to create high-quality and photorealistic images from simple textual prompts. Nevertheless, similar to other generative models (Mehrabi et al. 2021; Zhou et al. 2024), the widespread adoption of T2I models has raised significant concerns regarding fairness and biases, particularly in images depicting human (Bansal et al. 2022; Luccioni et al. 2023). Existing research highlights systematic biases within T2I models to specific social groups. For instance, T2I models primarily depict respectable occupations like lawyers and CEOs as white middle-aged men (Cho, Zala, and Bansal 2023), ignoring women and colored races.
Some researchers have proposed their solutions to address the issue (Gandikota et al. 2024; Schramowski et al. 2023; Shrestha et al. 2024). However, none of these methods fully resolve the two most crucial challenges that hinder the practical application of debiasing techniques. First, existing approaches are typically tailored to specific T2I models, commonly SD 1.5 (Rombach et al. 2022), which limits their adaptability, especially given the rapid evolution of T2I models driven by the adoption of distillation (Meng et al. 2023) and the participation of communities. For instance, the primary base model iterates three times from SD1.5 to SDXL-L (Lin, Wang, and Yang 2024) within a year, while researchers introducing models with entirely new architectures (Pernias et al. 2023; Chen et al. 2023). Recent research still develops their methods based on outdated models, limiting their application prospects. Second, most methods lack zero-shot debiasing capabilities. For example, FairDiffusion (Friedrich et al. 2023) uses a predefined look-up table to recognize prompts and add gender description to improve fairness, failing to deal with prompts that fall outside the predefined table. Like other AIGC techniques, effective debiasing techniques need robustness and zero-shot debiasing abilities to handle diverse inputs, ensuring broader applicability.
Figure 1: A simple generation workflow of VersusDebias.
To address this, we propose VersusDebias, a universal and self-adaptive zero-shot debiasing framework for T2I models, consisting of an array generation (AG) module and an image generation (IG) module. A simple workflow is shown in Figure 1. Referring the structure of generative adversarial networks (Goodfellow et al. 2020), our AG module employs a generator and a discriminator to perform generative adversary. The discriminator employs a multi-modal large language model (MLLM) to achieve exceptional alignment performance. The AG module produces debiased attribute arrays corresponding to 2,060 prompts, which consist of protected attributes from the ignored social groups, providing reference for modifying the prompts. Compared to existing methods that add protected attributes according to fixed demographic proportions, debiased attribute arrays mitigate the hallucinations of T2I models, i.e., instances where models generate images that are inconsistent with the prompt, through post-processing. The IG module employs a fine-tuned small language model (SLM), i.e., LLM with fewer parameters, which can accurately complete named entity recognition (NER) in complex prompts and contextually add protected attributes according to debiased attribute arrays to generate debiased prompts, providing zero-shot debiasing capability. Our SLM achieves LLM-level performance on NER tasks, surpassing traditional NER methods, while significantly reducing computational resource consumption compared to LLMs, thereby striking a outstanding balance. VersusDebias is able to debias any models without retraining, avoiding computational overhead. Moreover, as a universal plug-in module, VersusDebias is equipped with ComfyUI API (Comfyanonymous 2024) and Websocket, enabling it to adapt to thousands of models compatible with ComfyUI by simply replacing the workflow. This structure, with the adaptive capabilities provided by generative adversary, enables VersusDebias to debias any T2I models.
To evaluate the effectiveness of our work, we compare VersusDebias with other methods from two perspectives. First, we compare the demographic bias of three general models with and without VersusDebias. Second, we compare VersusDebias with existing debiasing methods. All comparisons consider both zero-shot and few-shot scenarios with multiple evaluation datasets. The results demonstrate that VersusDebias effectively reduces biases in T2I models and significantly outperforms other methods.
Our core contributions are summarized as follows:
  • We propose VersusDebias, a novel and universal framework to efficently decrease demographic biases in T2I models, which ensure fairness of the generated images.
  • VersusDebias integrates numerous advanced technologies, including MLLM, SLM, prompt engineering, and generative adversary.
  • VersusDebias implements zero-shot debiasing via the ComfyUI API and is transferable without retraining, showcasing its universality and promoting practical application of debiasing.
  • Experiments show that VersusDebias outperforms existing methods in both zero-shot and few-shot debiasing situations, proving its effectiveness and robustness.
2 Related Works
Biases in T2I Models
Over the last few years, T2I models like Stable Diffusion and DALL-E have seen increasing adoption (Ramesh et al. 2022; Esser et al. 2024). Many T2I models are able to generate high-quality images efficiently. Nonetheless, several studies prove that these models inherit biases from their training datasets (Chinchure et al. 2023; Wan et al. 2024). Recent research classifies these biases into implicit generative bias and explicit generative bias (Luo et al. 2024a), corresponding to the sociological concepts of implicit bias (Pritlove et al. 2019) and explicit bias (Fridell 2013). Implicit generative bias refers to the phenomenon where, without specific instructions on protected attributes, T2I models tend to generate images that do not consist of demographic facts. For instance, when asked to generate ”a photo of a surgeon”, models tend to produce images featuring male surgeons. Explicit generative bias is a specific type of hallucination, referring to the phenomenon where T2I models tend to generate images that do not consist with prompts containing specific protected attributes. For example, when asked to generate ”a photo of a rich black person” models may sometimes fail to correctly generate an image featuring a black person. In our work, we primarily address the issue of implicit generative bias and avoid the influence of explicit generative bias which disturbs existing methods.
Debiasing Based on Prompt Engineering
Recent works attempts to debias T2I models at different levels, such as diffusion-level (Shrestha et al. 2024) and prompt-level (Clemmer, Ding, and Feng 2024). Diffusion-level methods interfere with the image generation process, which can lead to uncontrollable side effects on the quality and style of the final image. Moreover, these methods can only be applied to diffusion-based models and are not suitable for transformer-based models like DALL-E3 (Betker et al. 2023). In contrast, debiasing based on prompt engineering, as a post-processing method, is more adjustable and more broadly applicable with less impact on image quality. These methods typically performs NER to identify the bias-inducing parts of the prompt and modifies them according to predefined proportions. However, these methods also face serious challenges. First, traditional prompt engineering methods rely on the assumption that the model correctly generate images based on the prompt, which highlights the problem of hallucinations, especially those caused by explicit generative bias. Second, prompt engineering might have unintended side effects on unrelated attributes. For example, using SDXL-Turbo (Sauer et al. 2023), the proportion of women generated with the prompt ”a photo of a South Asian tennis player” is significantly lower than that generated with the prompt ”a photo of a tennis player” (Luo et al. 2024b). Lastly, the severity of these issues varies across different models, meaning that traditional methods, even if they address the first two issues, can only be applied to specific models, greatly limiting their application. In contrast, our method employs an innovative AG module that addresses these problems, providing unprecedented practicality.
Figure 2: Pipeline of VersusDebias. The yellow boxes represent the components of the Discriminator. The three gray rectangles distinguish the three main stages. The blue rectangles represent data that change with the T2I model, while the red rectangles represent data that is independent of the T2I model.
Development of SLM and MLLM
Due to the tremendous success of ChatGPT-4 (Achiam et al. 2023), LLMs, especially MLLMs, garner significant attention and rapidly evolve. The improved accuracy of MLLMs addresses the significant shortcomings of traditional alignment models like CLIP (Radford et al. 2021) and BLIP-2 (Li et al. 2023) in recognizing race and age (Chinchure et al. 2023), providing essential support for the implementation of the alignment module in our discriminator. On the other hand, the development of edge AI attracts more researchers to focus on SLMs (Bai et al. 2023; Zhang et al. 2024), which retain most of the original performance of LLMs while significantly reducing computational resource consumption. Compared to traditional NER methods, LLM-based methods exhibit higher accuracy and zero-shot capabilities (Clemmer, Ding, and Feng 2024), which also indicates the potential of SLMs in NER. In VersusDebias, we utilize a recent SLM and a MLLM to achieve outstanding debiasing performance and accuracy with minimal computational resources.
3 Method
3.1 Overview
In this section, we introduce the pipeline of VersusDebias. A brief overview is shown in Figure 2. VersusDebias consists of a AG module, a IG module, and their datasets.
To debias a model, VersusDebias first uses prompts without protected attributes from the dataset to drive the generator to generate images, which are then conveyed into the discriminator. The discriminator employs a MLLM to perform semantic alignment on these images, extracting information on protected attributes. Due to the lack of statistic, we only consider the debiasing of gender, race, and age of the depicted persons. The alignment results are averaged to determine the generative demographic proportions of protected attributes for each prompt. The discriminator compares the proportions of each gender, race, and age group to the ground truth and modifies the prompts by adding the ignored protected attributes, which ensures that the generator produces more images featuring these ignored social groups in the next iteration. The added protected attributes are recorded in the protected attribute array corresponding to each prompt. The newly images generated from modified prompts are then conveyed into the discriminator again, and the process is repeated until the specified number of epochs.
The discriminator then evaluates the generation results for each prompt at each epoch, calculating the cosine similarity between the generative proportions and the ground truth to quantify the conformity between them. The protected attribute array with the highest cosine similarity for each prompt across different epochs is then combined into a single dictionary, referred to as the best results, and conveyed to the IG module.
For generating debiased images with the IG module, the executor uses the fine-tuned SLM to do NER accurately by extracting the potentially bias-inducing part from the input user prompt. It then selects an appropriate protected attribute based on the best results and adds it to the original prompt. Finally, the generator uses the processed prompt to generate debiased images.
3.2 Generator
As a universal debiasing framework based on prompt engineering, VersusDebias can use any model in its generator, including diffusion-based models (Nichol et al. 2021; Sauer et al. 2023; Esser et al. 2024; Lin, Wang, and Yang 2024; Song, Sun, and Yin 2024) and transformer-based models (Dayma et al. 2021; Ding et al. 2022; Ramesh et al. 2022; Chen et al. 2023). To lower the barrier of VersusDebias and attract more researchers to debiasing studies, we develop an interface with ComfyUI, which is a widely-used GUI for thousands of T2I models based on Stable Diffusion. ComfyUI only requires changes to the input workflow file to adapt to various models and is equipped with a comprehensive API for external program calls. Additionally, due to its openness and flexibility, some models that use new architectures, such as Playground (Li et al. 2024), PixArt (Chen et al. 2024b), and Stable Cascade, are also compatible with ComfyUI by simply installing specific extension nodes. In terms of communication, VersusDebias uses Websocket to monitor the status of ComfyUI and convey instructions. This approach not only achieves a fully automated debiasing process but also allows both systems to operate in independent environments, ensuring optimal compatibility.
In IG module, the generator uses the prompt processed by the executor to generate images based on user input. In AG module, the generator uses prompts from our dataset. The creation process for the dataset is as follows. Based on the 607 detailed occupations classified by the standard occupational classification of the U.S. (BLS 2024a), we consolidate occupations that are typically not distinguished in everyday contexts, resulting in 103 occupations. We also add descriptions of image quality and character states with ChatGPT-4o, generating 20 prompts for each occupation and totaling 2,060 prompts. The detailed process is provided in the supplementary material.
3.3 Discriminator
As the core part of the AG module, the discriminator consists of three components: the aligner, the array editor, and the evaluator. In the following sections, we provide a detailed explanation of the implementation of them.
Aligner
The aligner’s function is to align protected attributes with individuals in the images and compute the average of the responses of each occupation. We use the fine-tuned InternVL-4B-1.5 (Chen et al. 2024c) from BIGbench for alignment, which is trained with the dataset from FairFace (Karkkainen and Joo 2021) and reportedly achieves an accuracy of 97.93%. Using race alignment as an instance, the aligner asks the model ”Tell me the race of the main person in the image, White, Black, East Asian, South Asian, or unknown”, and store the response except ”unknown”. The aligner clean the history and tries again if ”unknown”, and skip the image if still ”unknown” persists, assuming the T2I model failed to generate human image. We do a test with results shown in the supplementary material, indicating that this model meeting the requirements of VersusDebias.
Array Editor
The function of the array editor is comparing the generative demographic proportions to the ground truth and adding the underrepresented protected attributes to the prompt. To achieve this, the array editor generates a new dictionary by subtracting the ground truth values from the actual generative proportions. The keys in this dictionary are the occupations, with negative values indicating that the generated data is below the actual data and positive values indicating the opposite. For all protected attributes with negative differences, the array editor adds these underrepresented attributes to the attribute array. The length of the attribute array $L$ can be adjusted based on the requirements for speed and accuracy, but it should generally be maintained at 4 times the number of epochs. This is because there are a total of nine protected attributes, and assuming an equal distribution of positive and negative differences, approximately 4.5 attributes are added each time. Allowing for a buffer, the multiplier is set to 4 as a floor value, and $L$ is at least 5 as a ceiling value. The array editor performs this operation once per epoch until the limit is reached. A detailed algorithm is shown in Algorithm 1.
Algorithm 1 Array Editing Algorithm
Input: Ground truth values $G$, actual generative proportions $A$, number of epochs $E$, length multiplier $m=4$
Output: Adjusted attribute array
1:Initialize dictionary $D\leftarrow\{\}$
2:for each prompt $p$ in $G$ do
3:     $D[p]\leftarrow A[p]-G[p]$
4:end for
5:Initialize attribute array $attrArray\leftarrow[]$
6:Set limit $limit\leftarrow 9\times E\times m$
7:for epoch $e$ from 1 to $E$ do
8:     for each attribute $a$ in $D$ do
9:         if $D[a]<0$ then
10:              Append $a$ to $attrArray$
11:         end if
12:     end for
13:     if length($attrArray$) $\geq$ $limit$ then
14:         break
15:     end if
16:end for
17:return $attrArray$
The ground truth uses the same occupational classification as the prompt set. For race, we employ the four-category classification from FairFace (Karkkainen and Joo 2021), distinguishing race into White, Black, East Asian, and South Asian. For age, we classify it into three stages: young (0-30), middle-aged (31-60), and elderly (60+), based on common age classifications. For gender, due to the cognitive limitations and the diverse of sexual minority identities, we classify gender into male and female, which limits our method to debias for sexual minorities. For race proportions, we use the global racial population data from the UN (UNDESA 2022), as VersusDebias is designed for worldwide users. For gender and age, the proportions for practitioners for a occupation are calculated by dividing the number of individuals of a specific group by the total number of individuals within all detailed occupation belonging to the occupation, based on the data from the BLS (BLS 2024b, c), due to a lack of global data on occupation practitioners and this data also following the SOC.
Evaluator
The evaluator provides a comprehensive metric for evaluating the overlap between the generated results and the ground truth when selecting the best-performing debiased attribute arrays. Since gender, race, and age each have 2-4 protected attributes, we treat them as vectors and use cosine similarity to determine the similarity between the generated results and the ground truth:
$S_{k}=\frac{\sum_{i=1}^{n_{k}}p_{ki}\cdot q_{ki}}{\sqrt{\sum_{i=1}^{n_{k}}p_{ki}^{2}}\cdot\sqrt{\sum_{i=1}^{n_{k}}q_{ki}^{2}}},$ (1)
where $S_{k}$ is the cosine similarity for the protected attribute $k$ of a occupation, $p_{k}i$ and $q_{k}i$ are the generative proportion and real demographic proportion of the $i$ attribute, and $n_{k}$ is the total number of the attributes for this protected attribute.
After calculating the cosine similarity for these three attributes separately, the evaluator computes a weighted average to obtain the cumulative cosine similarity. We set the weights for gender and race at 0.35 and the weight for age at 0.3, which reflects the general consensus that gender and racial discrimination is more common and severe. The weights can be adjusted based on different requirements:
$S_{c}=\sum_{k\in\{\text{gender},\text{race},\text{age}\}}w_{k}\cdot S_{k},$ (2)
where $S_{c}$ is the cumulative cosine similarity for a occupation, $w_{k}$ is the weight for the protected attribute $k$.
Finally, the evaluator selects the best-performing debiased attribute arrays based on the cumulative cosine similarity and combines them into a single dataset called ”best results”.
3.4 Executor
Figure 3: The pipeline of the executor. The blue rectangles represents the prompt, the red box represents the list of all occupations included in the prompt set, and the green box represents the debiased attribute arrays for R3.
The pipeline of the executor is briefly shown in Figure 3. The main component of the executor is a fine-tuned SLM. We use the SLM to accurately perform the following tasks in sequence: first, do NER to analyze practical prompts and recognize the potentially biased part; second, determine whether the prompt already includes protected attributes; third, identify a occupation in the occupation list that has the closest meaning to response 1. The executor locates the debiased attribute arrays corresponding to response 3 from the best results, randomly selects a protected attribute from the generated debiased attribute array, and adds the selected protected attribute to the prompt.
Due to the relatively fixed and simple nature of the tasks performed, we use a SLM instead of traditional LLMs to accomplish the aforementioned three tasks. SLMs are characterized by their faster operational speed and less resource consumption. However, they are subject to limitations in scale, which can lead to hallucinatory issues and less precise knowledge memory. To ensure accuracy, we develop a specialized dataset to fine-tune the model, thereby enhancing its capability to execute these tasks, particularly in the retrieval of synonyms. The dataset $D_{SLM}$ we used to train the SLM consists of 500 dialogue sets, with each set containing three rounds of queries $q$ and expected responses $r$, which is formulated as follows:
$D_{SLM}=\{(q_{i,k},r_{i,k})\mid i\in[1,500],k\in[1,3]\}$ (3)
In the first round, $q_{1}$ asks the model to identify the potentially biased part in a given prompt. $r_{1}$ is ”none” or the identified element, thus training the model’s ability of extracting. To closely simulate practical application scenarios, we use Promptomania (Promptomania 2024) to generate one prompt for each of the 103 occupations in the prompt set. Additionally, we use the official US occupational classification method (United States and Budget 2018) and thesaurus (Merriam-Webster 2023) to collect synonyms and variations for the occupations, resulting in a total of 380 queries. We test these prompts with SDXL (Podell et al. 2023) and manually revise prompts that produced poor image results to improve their quality. To ensure the SLM does not misidentify potentially biased parts, we also include 120 queries containing prompts that solely describe landscapes and objects to train the model’s judgment capabilities. In the second round, $q_{2}$ asks the model to determine whether the prompt originally contains a protected attribute, with $r_{2}$ being either ”yes” or ”no”. We included 260 sets without protected attributes and 120 sets with protected attributes to comprehensively train the model. In the third round, $q_{3}$ asks the model to select a synonym from the occupation list for $r_{1}$ if it’s not ”none” while $r_{3}$ is the chosen synonym.
4 Experiment
In this section, we initially present the comparative experiments conducted in the selection of SLMs. Subsequently, we summarize the performance of VersusDebias in debiasing various T2I models for both few-shot and zero-shot scenarios and investigate the influence of the attribute array length $L$. Finally, we compare VersusDebias with several baselines.
4.1 SLM Comparison
We select three state-of-the-art SLMs, Qwen2-1.5B (Yang et al. 2024), MiniCPM-2B (Hu et al. 2024), and Gemma-2B (Team et al. 2024). For comparative analysis, we choose representative LLMs deployable on a single GPU, Llama2-7B (Touvron et al. 2023) and Qwen2-7B, and the representative commercial large model, ChatGPT-4o (Achiam et al. 2023). All models except ChatGPT-4o are fine-tuned using our dataset $D_{SLM}$.
We conduct the analysis by developing a test set comprising 100 dialogue sets, consistent with the format of the training set, which consists of 52 dialogues describing humans without protected attributes, 24 describing humans with protected attributes, and 24 not involving human descriptions. With the test set, we evaluate the five models and calculated the accuracy by comparing the responses to the results of one human annotator undergoing specific training to evaluate the dialogues and occupations involved. We provide the datasheet in our respiratory for verification. Particularly, if $r_{1}$ can be categorized under multiple occupations in the element list, all of the occupations are correct for $r_{3}$. For instance, if $r_{1}$ is ”robotics engineer,” then $r_{3}$ is correct if its value is ”electrical engineer,” ”mechanical engineer,” or ”computer programmer”. The performance of models is detailed in Table 1. Qwen2-1.5B, while being the fastest in terms of speed with the usage of 5.4GB VRAM, achieves an accuracy that is slightly better than Llama2-7B and significantly higher than the other two SLMs, making it our choice. Furthermore, the performance of fine-tuned Qwen2-1.5B and Llama2-7B approaches ChatGPT-4o, which attests to the validity of our dataset design.
Model $r_{1}$ $r_{1}$-w/o $r_{2}$ $r_{3}$ Speed
ChatGPT-4o 100 100 100 98.65 -
Qwen2-7B 100 100 98.65 98.65 2.05
Llama2-7B 94.59 96.15 98.65 97.3 1.93
MiniCPM-2B 90.54 88.46 95.95 91.89 2.67
Gemma-2B 82.43 92.31 95.95 83.78 2.24
Qwen2-1.5B 97.3 100 98.65 95.95 4.23
Table 1: The $r_{1}$ column reflects the accuracy of the models when processing prompts that describe humans, while the $r_{1}$-w/o column reflects the accuracy for prompts without human descriptions. The Speed column indicates the average number of prompts processed per second with a RTX3090.
4.2 Debiasing Performance
Few-shot
For few-shot scenarios, we utilize ChatGPT-4o to generate 5 prompts for each of the 103 professions in the dataset and each prompt is used to generate 8 images. The format of these prompts imitates the pratical prompts. We employ these prompts to generate images and calculated the cosine similarity against the ground truth in the dataset. We calculate the cosine similarity between the results and ground truth of three T2I models, both with and without VersusDebias at a $L$ of 20. The two diffusion-based models are Stable Diffusion 1.5 (Rombach et al. 2022) and Stable Diffusion XL (Podell et al. 2023), while the transformer-based model is Pixart-$\Sigma$ (Chen et al. 2024a). These models are abbreviated as SDv1, SDXL, and PixArt in the following part. The result is shown in Table 2. The results fully demonstrate VersusDebias’s capability to effectively debias models of various principles in few-shot scenarios, especially in race and age where the models originally performed poorly. It is notable that PixArt, which initially performed worse, outperforms the originally better-performing SDv1 after debiasing. Upon reviewing the generated images, we find that this is because PixArt has a better capability to follow the guidance of the prompt. This highlights the importance of this capability for prompt-based debiasing.
Model Gender Race Age Total
SDv1 93.71 70.77 60.19 75.63
SDv1-VD 95.23 86.75 81.65 88.19
SDXL 89.15 65.52 78.85 77.79
SDXL-VD 96.24 85.29 91.44 90.97
PixArt 86.42 60.25 72.65 73.13
PixArt-VD 92.83 83.26 92.75 89.46
Table 2: Qualitative results in few-shot scenarios.
Zero-shot
For zero-shot scenarios, we design the evaluation based on BIGbench (Luo et al. 2024b). Firstly, we exclude the prompts from BIGbench if the occupations in the prompts overlap with our prompt set, and use the remaining 122 occupations and 2440 prompts to generate 8 images for each prompt. Subsequently, we calculate the cosine similarity. Like the evaluation of few-shot debiasing capabilities, we also conducted comparative experiments for SDv1, SDXL, and PixArt with and without VersusDebias. The result is shown in Table 3. For zero-shot scenarios, VersusDebias also achieves outstanding performance.
Model Gender Race Age Total
SDv1 91.36 68.13 53.94 72.00
SDv1-VD 95.12 87.59 77.21 87.11
SDXL 88.06 67.53 74.18 76.71
SDXL-VD 93.50 88.38 87.23 89.83
PixArt 85.92 58.87 70.25 71.75
PixArt-VD 90.73 86.16 87.21 88.07
Table 3: Qualitative results in zero-shot scenarios.
Model BIGbench total VersusDebias total PreciseDebias total
gender race age gender race age gender race age
SDv1 91.36 68.13 53.94 72.00 93.71 70.77 60.19 75.63 85.34 58.90 51.78 66.02
FairDiffusion 95.90 73.80 64.14 78.64 95.12 69.18 62.23 76.17 92.19 61.81 52.63 69.69
PreciseDebias 86.92 95.56 54.14 80.11 84.55 96.17 58.21 80.72 77.83 80.63 58.79 73.10
VersusDebias 95.12 87.59 73.21 85.91 95.23 86.75 81.65 88.19 91.69 72.73 75.41 80.17
Table 4: VersusDebias exhibits the best overall performance and balanced results across attributes, demonstrating its comprehensiveness. We provide more detailed experimental results and further discuss the differences in the supplementary material.
Influence of Array Length $L$
Due to VersusDebias’s reliance on randomly selecting protected attributes from debiased attribute arrays, the probability of selecting a protected attribute is quantized. Therefore, a larger attribute array can result in smoother variations in generative demographic proportions, improving accuracy. As explained in Section 3.3, the length of the debiased attribute array $L$ is greater than or equal to 5. We conduct comparative experiments with $L$ set to 5, 10, 20, 50, and 100, corresponding to the epochs being 2, 3, 5, 13, and 25, respectively. The T2I model used in the experiment is SDv1. The result is shown in Figure 4.
Figure 4: Performance curve with the increase of $L$.
Figure 5: With VersusDebias, the generated results encompass all genders, races, and ages, and proportions closely matches the ground truth.
The x-axis of the graph represents $L$, and the y-axis represents cosine similarity in the best results for few-shot or the results of BIGbench for zero-shot. An array size ”0” indicates the performance without VersusDebias. The results show that larger array lengths effectively enhance the performance, with a clear marginal effect. Therefore, we set $L$ to 20 in other experiments to achieve a balance between performance and speed. Detailed computational resource consumption is provided in the supplementary material.
4.3 Comparison with Baselines
As no prior universal debiasing methods have been developed, we select two prompt-level existing methods based on SDv1, FairDiffusion (Friedrich et al. 2023) and PreciseDebias (Clemmer, Ding, and Feng 2024). We also use SDv1 for VersusDebias in this part of the experiment to control variables. FairDiffusion uses a look-up table to do NER and adds gender descriptions based on predefined probabilities for different occupations. PreciseDebias employs a fine-tuned Llama2-7B model to do NER and adds gender and racial descriptions according to fixed general probabilities. In the contrast, VersusDebias debiases gender, age, and race simultaneously via the AG module without knowing all specific proportions. To ensure the fairness, we replace the original racial proportions in PreciseDebias with the proportions from the ground truth and include additional test based on the dataset of PreciseDebias. Table 4 shows the results and Figure 5 provides a visualized example.
5 Limitation and Future Work
Although VersusDebias significantly outperforms existing debiasing methods, our framework still need to be improve in some aspects. First, although the fine-tuned InternVL-4B-1.5 demonstrates excellent alignment, particularly in racial alignment, it still cannot achieve perfect alignment. This limitation reduces the accuracy of the discriminator, which in turn affects the performance of the discriminator. A better model that is specifically designed for human classification is worth investigating. Secondly, as a post-processing method, VersusDebias fails to handle extreme cases of hallucinations, such as when the model completely fails to generate images that match the prompt ”an Asian husband and a white wife” (Vincent 2024). This issue needs model-level improvement. Lastly, since VersusDebias only solves the problem of implicit bias, explicit bias remains to be addressed in future work. We hope our framework can be helpful for future research about explicit bias.
6 Conclusion
This paper introduces VersusDebias, a universal debiasing framework for T2I models, which enables precise and efficient debiasing for any T2I models by prompt engineering. To achieve this, we design a novel AG module to generate debiased attribute arrays for specific models and apply a advanced fine-tuned SLM to do NER and edit prompts according to the arrays. The edited prompts are used in T2I models to generate images with less demographic biases. Our AG module addresses the issues of hallucinations and side effects caused by prompt engineering, which are not handled by previous prompt-level methods, significantly improving the performance of our method. We also use ComfyUI API to lower the barriers of debiasing. VersusDebias surpasses existing methods in both zero-shot and few-shot scenarios, while maintaining the models’ image style and quality. We hope that VersusDebias can promote the application of debiasing methods in T2I models, for a fairer AIGC community.